Activity has been at a relatively low level during this reporting period. The October 22nd dome continues to grow on the southern side while the northern flanks remain quiet. Seismicity has been dominated by rockfall signals with two periods of pyroclastic flow generation on the 27th November and 1st December.
During the last week visibility has significantly improved and good observations were made of the dome. Dome growth continues to be concentrated on the southern side of the dome in the Galway's area. A small dome collapse occurred on the 27th November at 14:45 and continued until after 18:00. Although some of these flows reached the fan, the majority of the collapse was made up of continuous but relatively small flows reaching less than 2km.
There was further pyroclastic flow activity on 1st December which started around 8 pm. This was preceded by 2 hours of elevated rockfall activity. Again only the largest of these pulses reached as far as the delta. While the flows were occurring glowing was observed in the sky above the Galway's area. Neither the flows on the 27th nor those on the 1st were considered to have produced ash clouds above 10,000 ft, or to have involved significant volumes of material.
Good views of the dome were obtained on 4th December for the first time since 26th November. During this time the talus apron at the base of the dome in the Galway's area had significantly built up both in height and width. The continued growth of the October 22nd dome in this area is beginning to encroach on Chances Peak and also on the east in the Tar River side. The height of the active centre on the southern side is well above the height of the explosion crater rim and is probably around 980m. Theodolite surveys have confirmed that the northern lobe of the October 22nd dome has not grown during this period. Occasional minor rockfalls have occurred on the northern and eastern face in the Tar River valley area as old surfaces of the dome gradually degrade.
The reporting period is characterised by a further decrease of seismic activity that had begun around 14th November. Most remarkable is the lack of hybrid earthquakes in the second week while the number of rockfall events remained at a constant level. The frequency contents of the rockfall signals changed slightly towards longer periods probably due to the attenuation of higher frequencies. This points to a possible change of ground coupling between rock falls or pyroclastic flows and unconsolidated deposits on the talus slope.
Table 1: Earthquake types
These earthquake counts are of events that triggered the broadband network's event recording system between 16:00 and 16:00 each day (local time).
Date VT Hybrid LP Dome RF LPRF* HYRF* 23 Nov 97 0 74 19 44 1 1 24 Nov 97 0 102 8 37 1 0 25 Nov 97 0 16 24 57 4 0 26 Nov 97 0 9 5 27 5 0 27 Nov 97 0 17 8 84 0 5 28 Nov 97 1 11 17 71 1 1 29 Nov 97 0 7 7 68 5 0 30 Nov 97 0 0 12 50 3 0 1 Dec 97 0 9 26 79 3 1 2 Dec 97 1 8 24 73 12 4 3 Dec 97 0 8 30 92 15 8 4 Dec 97 2 7 30 91 5 1 5 Dec 97 0 5 22 128 3 0 6 Dec 97 0 9 35 97 3 8 7 Dec 97 0 7 35 121 11 5
* LPRF: LP earthquake followed by rockfall signal. HYRF: Hybrid earthquake followed by rockfall signal. The LPs, hybrids and rockfalls in these signals are also counted in their respective columns.
Occupations of the GPS networks BIGNET (Harris, Whites, Windy Hill and Broderick's) and LEESNET (Old Towne, WaterWorks, St George's Hill and Lees Yard) were made during this reporting interval. The baseline Harris-Whites appears to have recovered half of the shortening that has occurred over the last five months. The line is now only around 2cm shorter than its pre-June mean. The data suggest a slow rise of the Broderick's site, although this is not yet certain. No clear trends have yet been identified in the data collected on LEESNET.
A deposit survey was undertaken on 28th November. This covered all the area between Mosquito Ghaut in the north and White's Ghaut in the east including the Farm, Trant's and the airport area. The total volume increase of material in these ghauts since 13th August is 24 million cubic metres. This volume is partitioned between the collapse of the 21st September, the column collapse flows produced during the explosive activity and pyroclastic flow activity on the northern flank between mid August and the 21st September event.
The total volume of material erupted now stands at 203 million cubic metres, this includes the dome itself, the pyroclastic flow material and estimates for the eruption column material. The previous dome survey gave an extrusion rate of 5m3/s which is a decrease from the summer months but the overall trend for the average extrusion rate is still increasing.
Dust Trak sampling has been carried out at four fixed sites. Each value is an average of the PM10 concentration over approximately 24 hrs. The values at the fixed sites have been low over this period except for the values at the Catholic school which sometimes recorded raised levels, this is due to the large amount of human activity at this site and the fact that it is near a main road.
On the 6th and 7th of December the Dust Traks were taken in to be downloaded and for some experiments to be carried out.
Table 3 : PM10 (concentration in mg/m3)
These values are approximate 24 hr averages of the PM10 concentration.
Site 24/11/97 25/11/97 26/11/97 27/11/97 28/11/97 29/11/97 30/11/97 Mango 0.017 0.012 0.019 0.018 0.01 0.015 0.019 St. Peter's 0.017 0.01 0.013 0.01 0.011 0.018 0.02 CPS 0.034 0.031 0.04 0.024 0.024 0.043 0.021 MVO north 0.020 0.013 0.012 0.01 0.008 0.014 0.02 Site 01/12/97 02/12/97 03/12/97 04/12/97 05/12/97 06/12/97 07/12/97 Mango 0.011 0.011 0.01 0.009 0.008 - - St. Peter's 0.01 0.011 0.01 0.01 0.009 - - CPS 0.031 0.045 0.044 0.056 0.063 - - MVO north 0.011 0.01 0.008 0.007 0.007 - -
24 Hour averages: <0.05 mg m-3 Low 0.05-0.1 mg m-3 Raised 0.1-0.3 mg m-3 Very High >0.3 mg m-3 Alert
St. Peter's: The Dust Trak is outside on the balcony of a villa in St Peter's, called St. Peter's Place. CPS: Catholic Primary School The Dust Trak is outside in the area under the roof where some of the children have classes. The school is in Palm Loop near Woodlands. MVO north: The new Volcano observatory on Mongo Hill near St.John's, the Dust trak is on the third story of the building with the sampling tube stuck out of the window. Mango: The Dust Trak is outside on the verandah of a villa in Mango Drive in Woodlands
MVO Staff Changes
Jurgen Neuberg (Leeds University, U.K.)
Caroline Choux (Clermont-Ferand University, France)
Richard Luckett (MVO) left for a short visit to the States.
Rob Watts (Independent/University of Bristol)
On the 3rd October, a contingent of senior scientists from the Antigua scientific summit meeting.